Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 260     Entry time: 18 Jun 2023, 07:31
Author: Adam Round 
Group: SPB 
Subject: Shift end 

posting for 17th June

Staff: AR, AS, FK, HH, JB, KK, DZ, MK

X-ray delivery

  • 12.4 keV, 1.4 mJ

Optical laser delivery

  • 532 nm, ns-laser for sample pumping

Achievements/Observations

  • Continued to collect data many hours of data collection

Issues

  • coupling with SASE2 limited feedbacks to control slope across train, this led to dificulties in alignment and lost time to correct for poor visualisation of the beam.
  • After discussion SASE2 kindly gave permission to BKR to corect the slope to facilitate our alignment
    • we also with the help of PRC arranged for the same arrangment for alignment for sunday
ELOG V3.1.4-7c3fd00